OpenStudy (anonymous):

find the x-intercept of the parabola with vertex (-3,-14) and y intercept (0,13) in (x1,y1) (x2,y2) form using - (y-k)=a(x-h)^2

OpenStudy (anonymous):

vertex is (-3,-14) means you know it looks like \[y=a(x+3)^2-14\]

OpenStudy (anonymous):

and since you know that if x = 0, y = 13 replace x by 0 and set y = 13 to get a \[13=a(0+3)^2-14\] \[13=9a-14\] \[27=9a\] \[a=3\]

OpenStudy (anonymous):

thus \[y=3(x+3)^2-14\] or if you prefer \[y+14=3(x+3)^2\]

OpenStudy (anonymous):

so thats what i type? it saids in (x1,y1) (x2,y2) form ?

OpenStudy (anonymous):

i have no idea what that means. it looks like you want \[y+14=3(x+3)^2\]